# Demystifying Deep Reinforcement Learning-Based Autonomous Vehicle Decision-Making

### v1

@@ Abstract
With the advent of universal function approximators in the domain of reinforcement learning, the number of practical applications leveraging deep reinforcement learning (DRL) has exploded. Decision-making in automated driving tasks has emerged as a chief application among them, taking the sensor data or the higher-order kinematic variables as the input and providing a discrete choice or continuous control output. However, the black-box nature of the models presents an overwhelming limitation that restricts the real-world deployment of DRL in autonomous vehicles (AVs). Therefore, in this research work, we focus on the interpretability of an attention-based DRL framework. We use a continuous proximal policy optimization-based DRL algorithm as the baseline model and add a multi-head attention framework in an open-source AV simulation environment. We provide some analytical techniques for discussing the interpretability of the trained models in terms of explainability and causality for spatial and temporal correlations. We show that the weights in the first head encode the positions of the neighboring vehicles while the second head focuses on the leader vehicle exclusively. Also, the ego vehicle’s action is causally dependent on the vehicles in the target lane spatially and temporally. Through these findings, we reliably show that these techniques can help practitioners decipher the results of the DRL algorithms.
@@ I Introduction
The answer lies in the black-box nature of the DRL and the difficulty that arises in understanding the intention behind the action. For a safety-critical application such as AVs, a sub-optimal but explainable solution is clearly preferred. In this paper, we hope to provide some scientific insights into that black box using some interpretation tools. Specifically, we focus on the attention-based mechanism which provides weights on the states (in this case, the neighboring vehicles) to interpret the policies. Our novel contributions are as follows:
@@ II-A AV Decision-making
Two kinds of methods are generally utilized for AV decision-making, classical and learning-based methods [9]. Classical methods, such as FSM [10, 11], Model Predictive Control (MPC) [12, 13, 14], game theory, etc., have been widely used for AV decision-making tasks including path tracking, car-following, and lane-changing. The advantages of classical methods are their simplicity and strong interpretability. Moreover, methods such as MPC and game theory are able to model the interactions between different traffic participants. However, rule-based methods in general have poor performance in complex driving scenarios as their parameters are usually hand-tuned after simulation and field testing [15].
Learning-based methods, in particular, deep learning, offer several benefits for vehicle control compared to classical methods. Deep learning methods have shown great promise in not only providing excellent performance for complex and non-linear control problems but also in generalizing previously learned rules to new scenarios [15]. Among existing deep-learning methods, model-free DRL methods have become more and more popular for various AV control tasks, including lane-change, car-following, path-tracking, and platooning [16, 17, 18, 19]. DRL methods treat AVs as agents, which interact with the environments and learn the consequences of their actions through experience via algorithms such as policy gradient, Q-Learning, etc. As they are model-free, they do not require specific models to measure the environments, and the agents can learn directly from experience. This makes them well-suited for environments where the dynamics are unknown or complicated to model, such as various driving scenarios.
@@ II-B Explainability and Causality in AV Decision-making
In addition to achieving safer and more efficient AV decision-making, explainability and causality are two crucial considerations for evaluating AV decision-making systems, especially for deep learning methods. The need for them is due to the increasing concerns for transparency and accountability of AVs [20]. Specifically, explainability refers to the ability to explain and justify the underlying rationale of AVs’ decisions. Differently, causality explores the cause-and-effect relationships between variables, such as the actions of neighbor vehicles and the actions of AVs. It aims to determine whether changes in one variable directly lead to changes in another variable, rather than just identifying similarities. Although DRL-based models have made significant progress and achieved superior performance in AV control compared to classical models, one major limitation of DRL-based models is the lack of explainability and causality. To solve this problem, explainable reinforcement learning, known as XRL, has been proposed and has gained a lot of attention in recent years. The main focus of XRL is on the reasoning behind the decisions or predictions made by the RL model [21]. Specifically, the majority of XRL methods function by mimicking and simplifying a complex model using a more explainable model, such as tree-based models [22, 23, 24]. Specifically, a DRL policy is first trained for controlling the AV. Then, a decision tree model is developed to explain the model using the observations of the agent as inputs and the decisions generated by the policy as outputs. This approach is easy to implement and interpret, however, the explainability may not reveal the actual knowledge learned by the agent. To further improve the explainability of DRL-based AV control models, additional research needs to be done on improving the inherent explainability of the DRL model rather than using a simple model to imitate it. Moreover, little research has been done to understand the causality behind deep learning-based methods for AV decision-making. [25] attempts to utilize a causality test for understanding the causality of a vehicle’s actions using image data. A visual attention model is trained end-to-end from images to the steering angle. The causality is then decided by iteratively masking attention regions in images and evaluating the impact on the model’s performance. This casual filtering technique produces a simpler and more accurate map of visual saliency by filtering out spurious attention regions. However, causality was not further analyzed but served as a technique to improve the interpretation of the attention maps.
@@ III-D1 Reinforcement learning
Reinforcement learning is a general probabilistic framework for sequential decision-making under uncertainty. Based on the feature we extracted, we train an action predictor in reinforcement learning using the proximal policy optimization (PPO) algorithm [32]. We train our agent in highway-env using PPO as the reinforcement learning algorithm that takes the $feat$ from the feature extractor as input and outputs an action $a\in\mathbb{R}^{2}$.
PPO is an RL architecture that improves training stability by avoiding large policy updates. The intuition is that performing smaller policy updates during training converges to an optimal solution more often. Therefore, the change is clipped to $\epsilon$ over the current policy. Its loss function is
@@ IV-A Benchmarking DRL agent performance
We train the baseline model and the 1-, 2- and 4-head attention models over the 3M global steps. Figure 2 shows the reward over the iterations for the different heads. We find that the 2-head fares better compared to others in collecting rewards over the course of the training.
@@
Figure 4: Change in attention weights and distance during left lane change and right lane change.
@@ IV-B2 Temporal explainability of AV DRL
We investigate the temporal explainability by plotting the changes in attention weights when the AV performs a left lane change from the right lane to the middle lane as shown in Figure 5111The full video including the weights are given in https://drive.google.com/file/d/1niiYw1GgUFWJcuzsVfSiKKdABtB6aZkb/view?usp=drive_link. The weights of the left leader from the two attention heads gradually increase as the AV starts the preparation of the lane change. They then reach the maximum values when the AV performs the lane change. This result corresponds to the reality as the AV would pay more attention to the left leader since it needs to find the appropriate gap to change the lane. On the contrary, the weights of the leader gradually decrease as the AV starts performing the lane change. The AV pays more attention to the leader in the beginning as it needs to maintain a safe distance to the leader. However, it focuses less on the leader while performing the lane change as this action does not affect the leader directly.

### v2

# Demystifying the Physics of Deep Reinforcement Learning-Based Autonomous Vehicle Decision-Making

@@ Abstract
With the advent of universal function approximators in the domain of reinforcement learning, the number of practical applications leveraging deep reinforcement learning (DRL) has exploded. Decision-making in autonomous vehicles (AVs) has emerged as a chief application among them, taking the sensor data or the higher-order kinematic variables as the input and providing a discrete choice or continuous control output. There has been a continuous effort to understand the black-box nature of the DRL models, but so far, there hasn’t been any discussion (to the best of authors’ knowledge) about how the models learn the physical process. This presents an overwhelming limitation that restricts the real-world deployment of DRL in AVs. Therefore, in this research work, we try to decode the knowledge learnt by the attention-based DRL framework about the physical process. We use a continuous proximal policy optimization-based DRL algorithm as the baseline model and add a multi-head attention framework in an open-source AV simulation environment. We provide some analytical techniques for discussing the interpretability of the trained models in terms of explainability and causality for spatial and temporal correlations. We show that the weights in the first head encode the positions of the neighboring vehicles while the second head focuses on the leader vehicle exclusively. Also, the ego vehicle’s action is causally dependent on the vehicles in the target lane spatially and temporally. Through these findings, we reliably show that these techniques can help practitioners decipher the results of the DRL algorithms.
@@ I Introduction
The answer lies in the black-box nature of the DRL and the difficulty that arises in understanding the intention (based on physics) behind the action. For a safety-critical application such as AVs, a sub-optimal but explainable solution is clearly preferred. In this paper, we hope to provide some scientific insights into that black box using some interpretation tools. Different from other papers in the explainability domain, we focus on deciphering the physics that the model learns, which is important for the downstream engineering tasks. Specifically, we concentrate on the attention-based mechanism which provides weights on the states (in this case, the neighboring vehicles) to interpret the policies. Our novel contributions are as follows:
@@ II-A AV Decision-making
AV decision-making can be broken down into two different types - classical and learning-based methods [9]. Classical methods, such as FSM [10, 11], Model Predictive Control (MPC) [12, 13, 14], and game theory [15], etc., have been widely used for AV decision-making tasks including path tracking, car-following, and lane-changing. The classical methods offer significant advantages in terms of simplicity and strong interpretability. Moreover, methods such as MPC and game theory can model the interactions between different traffic participants. However, classical methods may have poor performance in complex scenarios as their parameters are usually hand-tuned after simulation and field testing [16].
Learning-based methods, particularly deep learning, offer several advantages for vehicle control over classical methods. Deep learning methods excel in handling complex and non-linear control and can generalize previously learned rules to new scenarios [16]. Two common deep learning methods, including supervised learning and reinforcement learning, are commonly used in existing studies. Supervised learning utilizes labeled data, where an expert demonstrates the task for training purposes [16]. The labeled data contains observation-action pairs, such as images from onboard cameras and corresponding steering angle values. During training, the model learns how to act based on the observations and the actions. Supervised learning is easy and efficient to implement. However, it may not generalize well as it is heavily dependent on the training data. Differently, reinforcement learning allows the model to learn actions through trial and error and does not require labeled data. Therefore, models developed using reinforcement learning can generalize well to new scenarios. Among existing deep-learning methods, model-free DRL methods are becoming more and more popular as they allow adaptive and flexible learning from raw experiences without needing a predefined model, enhancing real-time performance in dynamic environments. They have been used in various AV control tasks, including lane-change [17, 18], car-following [19], path-tracking, and platooning [20]. DRL methods treat AVs as agents, which interact with the environments and learn the consequences of their actions through experience via algorithms such as policy gradient, Q-Learning, etc. As they are model-free, they do not require specific models to measure the environments, and the agents can learn directly from experience. This makes them well-suited for complicated driving environments where the dynamics are unknown or complicated to model.
@@ II-B Explainability and Causality in AV Decision-making
Although deep learning models have achieved promising accuracy in AV decision-making. There have been concerns over their transparency and accountability [21] since they are known as ”black-box” models. Therefore, explainability and causality have become necessary criteria while evaluating deep learning-based AV decision-making systems. Specifically, explainability refers to the ability to explain and justify the underlying rationale of AVs’ decisions. Differently, causality explores the cause-and-effect relationships between variables, such as the actions of neighbor vehicles and the actions of AVs. It aims to determine whether changes in one variable directly lead to changes in another variable, rather than just identifying similarities. Although DRL-based models have made significant progress and achieved superior performance in AV control compared to classical models, one major limitation of DRL-based models is the lack of explainability and causality. To solve this problem, explainable reinforcement learning, known as XRL, has been proposed and has gained a lot of attention in recent years. The main focus of XRL is on the reasoning behind the decisions or predictions made by the RL model [22]. Specifically, the majority of XRL methods function by mimicking and simplifying a complex model using a more explainable model, such as tree-based models [23, 24, 25]. Specifically, a DRL policy is first trained for controlling the AV. Then, a decision tree model is developed to explain the model using the observations of the agent as inputs and the decisions generated by the policy as outputs. This approach is easy to implement and interpret, however, the explainability may not reveal the actual knowledge learned by the agent. To further improve the explainability of DRL-based AV control models, additional research needs to be done on improving the inherent explainability of the DRL model rather than using a simple model to imitate it. Moreover, little research has been done to understand the causality behind deep learning-based methods for AV decision-making. [26] attempts to utilize a causality test for understanding the causality of a vehicle’s actions using image data. A visual attention model is trained end-to-end from images to the steering angle. The causality is then decided by iteratively masking attention regions in images and evaluating the impact on the model’s performance. This casual filtering technique produces a simpler and more accurate map of visual saliency by filtering out spurious attention regions. However, causality was not further analyzed but served as a technique to improve the interpretation of the attention maps.
@@ III-D1 Reinforcement learning
Based on the feature we extracted, we train an action predictor in reinforcement learning using the proximal policy optimization (PPO) algorithm [33]. We train our agent in highway-env using PPO as the reinforcement learning algorithm that takes the $feat$ from the feature extractor as input and outputs an action $a\in\mathbb{R}^{2}$. PPO is an RL architecture that improves training stability by avoiding large policy updates. The intuition is that performing smaller policy updates during training converges to an optimal solution more often. Therefore, the change is clipped to $\epsilon$ over the current policy. Its loss function is
@@ IV-A Benchmarking DRL agent performance
We train the baseline model and the 1-, 2- and 4-head attention models over the 3M global steps (3-head did not provide any results). Figure 2 shows the reward over the iterations for the different heads. We find that the 2-head fares better compared to others in collecting rewards over the course of the training.
@@
Figure 4: 2D histogram showing the change in attention weights with respect to normalized distance to the leader vehicle during cumulative left and right lane changes - the weights can be shown to have direct correlation with distance although the rate of change differs significantly
@@ IV-B2 Temporal explainability of AV DRL
We investigate the temporal explainability by plotting the changes in attention weights when the AV performs a left lane change from the right lane to the middle lane as shown in Figure 5111The full video including the weights are given in https://drive.google.com/file/d/1viFwTF_zxDWyRKnwOkUjKvtDHeey7gXz/view?usp=share_link. The weights of the left leader from the two attention heads gradually increase as the AV starts the preparation of the lane change. They then reach the maximum values when the AV performs the lane change. This result corresponds to the reality as the AV would pay more attention to the left leader since it needs to find the appropriate gap to change the lane. On the contrary, the weights of the leader gradually decrease as the AV starts performing the lane change. The AV pays more attention to the leader in the beginning as it needs to maintain a safe distance to the leader. However, it focuses less on the leader while performing the lane change as this action does not affect the leader directly.